one positive number is 1/5 of another number. the difference between the two numbers is 28. find the numbers
Answer by stanbon(75887)

one positive number is 1/5 of another number. the difference between the two numbers is 28. find the numbers
----
Equations:
x = (1/5)y
y - x = 28
------
Substitute for "x" and solve for "y":
y - (1/5)y = 28
(4/5)y = 28
y = (5/4)28
y = 35
-----
Solve for "x":
x = (1/5)y
x = (1/5)35
x = 7
=================
Cheers,
Stan H.